11 11 TEX L^{A}TEX

From DOS/Windows to Linux HOWTO

By Guido Gonzato, <REMOVE_MEguido@ibogeo.df.unibo.it> (Remove “REMOVE_ME”) v1.3.2, 22 February 1999

This HOWTO is dedicated to all the (soon to be former?) DOS and Windows users who have decided to switch to Linux, the free UNIX clone. The purpose of this document is to help the reader translate his or her knowledge of DOS and Windows into the Linux environment, as well as providing hints on exchanging files and resources between the two OSes.

Contents

1 Introduction 2 1.1 Is Linux Right for You? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 It Is. Tell Me More . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1.2.1 Introductory Concepts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2.2 Getting Help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

1.3 Conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 For the Impatient 5

3 Meet bash 5

4 Files and Programs 6 4.1 Files: Preliminary Notions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 4.2 Symbolic Links . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.3 Permissions and Ownership . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.4 Files: Translating Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

4.4.1 Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.5 Running Programs: Multitasking and Sessions . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.6 Running Programs on Remote Computers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

5 Using Directories 11 5.1 Directories: Preliminary Notions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 5.2 Directories Permissions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 5.3 Directories: Translating Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

5.3.1 Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

6 Floppies, Hard Disks, and the Like 13 6.1 Managing Devices the DOS Way . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 6.2 Managing Devices the UNIX Way . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 6.3 Backing Up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

7 What About Windows? 15

8 Tailoring the System 16 8.1 System Initialisation Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 8.2 Program Initialisation Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

9 Networking: Concepts 18

10 A Bit of Programming 18 10.1 Shell Scripts: .BAT Files on Steroids . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 10.2 C for Yourself . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

11 The Remaining 1% 21 11.1 Using tar and gzip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 11.2 Installing Applications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 11.3 Tips You Can’t Do Without . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 11.4 Where to Find Applications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 11.5 A Few Things You Couldn’t Do . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 11.6 Practicing UNIX under DOS/Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 11.7 Common Extensions and Related Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 11.8 Converting Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 11.9 Free Office Suites . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

12 The End, for Now 26 12.1 Copyright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 12.2 Disclaimer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

1 Introduction

1.1 Is Linux Right for You?

You want to switch from the DOS world (this definition includes Windows) to Linux? Good idea: Linux is technically superior to DOS, Windows 9x and even Windows NT. But beware: it might not be useful for you. These are the main differences between DOS/Windows and Linux:

• Windows runs Microsoft Office and lots of games; is perceived to be easy to install and configure; is notoriously unstable; performs poorly; crashes are frequent.

• Linux runs StarOffice, scores of technical software and fewer games; can be difficult to install and configure; is rock solid; performs impeccably; crashes are extremely rare.

It’s up to you to decide what you need. Furthermore, Linux gives you power, but it takes some time to learn how to harness it. Thus, if mostly need commercial sw, or if you don’t feel like learning new commands and concepts, you had better look elsewhere. Be aware that many newcomers give up because of initial difficulties. Work is underway to make Linux simpler to use, but don’t expect to be proficient with it unless you read a lot of documentation and use it at least for a few months. Linux won’t give you instant results. In spite of these warnings, I’m 100% confident that if you are the right user type you’ll find in Linux your computer Nirvana. By the way, Linux + DOS/Win can coexist happily on the same machine. Prerequisites for this howto: I’ll assume that

• you know the basic DOS commands and concepts;

• Linux, possibly with X Window System (X11 for short), is properly installed on your PC;

• your shell (the equivalent of COMMAND.COM) is bash.

Unless specified, all information in this work is aimed at bad ol’ DOS. There is information about Windows here and there, but bear in mind that Windows and Linux are totally different, unlike DOS that is sort of a UNIX poor relation.

Please also note that this work is neither a complete primer nor a configuration guide!

The latest version of this document is available in several formats on <ftp://sunsite.unc.edu:/pub/

Linux/docs/howto>.

1.2 It Is. Tell Me More

You installed Linux and the programs you needed on the PC. You gave yourself an account (if not, type adduser yourname now!) and Linux is running. You’ve just entered your name and password, and now you are looking at the screen thinking: “Well, now what?”

Now, don’t despair. You’re almost ready to do the same things you used to do with DOS/Win, and many more. If you were running DOS/Win instead of Linux, you would be doing some of the following tasks:

• running programs and creating, copying, viewing, deleting, printing, renaming files;

• CD’ing, MD’ing, RD’ing, and DIR’ring your directories;

• formatting floppies and copying files from/to them;

• tailoring the system;

• surfing the Internet;

• writing .BAT files and programs in your favourite language;

• the remaining 1%.

You’ll be glad to know that these tasks can be accomplished under Linux in a fashion similar to DOS. Under DOS, the average user uses very few of the 100+ commands available: the same, up to a point, applies to Linux.

1.2.1 Introductory Concepts

The best way to learn something new is to get your feet wet. You are strongly encouraged to experiment and play with Linux: unless you login as “root”, you can’t damage the system that way. A few points:

• first of all, how to quit Linux safely. If you see a text mode screen, press <CTRL-ALT-DEL>, wait for the system to reboot, then switch off the PC. If you are working under X Window System, press <CTRL-ALT-BACKSPACE> first, then <CTRL-ALT-DEL>. Never switch off or reset the PC directly: this could damage the file system;

• unlike DOS or Windows, Linux has built-in security mechanisms. Files and directories have permissions associated to them; as a result, some cannot be accessed by the normal user; (see Section 4.3 (Permissions and Ownership)). DOS and Windows, on the contrary, will let you wipe out the entire contents of your hard disk; • there’s a special user called “root”: the system administrator, with full power of life and death on the

machine. If you work on your own PC, you’ll be root as well. Working as root is dangerous: any mistake can seriously damage or destroy the system just like with DOS/Win. Don’t work as root unless absolutely necessary;

• much of the complexity of Linux comes from its extreme configurability: virtually every feature and every application can be tailored through one or more configuration files. Complexity is the price to pay for power;

• redirection and piping are a side DOS feature, a very inportant one and much more powerful under Linux. Simple commands can be strung together to accomplish complex tasks. I strongly suggest that you learn how to use them.

1.2.2 Getting Help

There are many ways to get help with Linux. The most important are:

• reading the documentation—I mean it. Although the HOWTO you are reading may serve as an introduction to Linux, there are several books that you really should read: Matt Welsh’s “Linux Installation

and Getting Started” ( <http://sunsite.unc.edu/mdw/LDP/gs/gs.html>), Larry Greenfield’s “Linux
User Guide” ( <ftp://sunsite.unc.edu/pub/Linux/docs/linux-doc-project/users-guide>), and
the Linux FAQ ( <http://sunsite.unc.edu/mdw/FAQ/Linux-FAQ.html>). Feel a guilty conscience

until you have read at least one of them;

• the documentation of the packages installed on the machine is often found in subdirectories under

/usr/doc/;

• to get some help about the “internal commands” of the shell, type help or, better, man bash or info

bash;

• to get help about a command, type man command that invokes the manual (“man”) page of command. Alternatively, type info command that invokes, if available, the info page pertinent of command; info is a hypertext-based documentation system, perhaps not intuitive to use at first. Finally, you may try apropos command or whatis command. With all of these commands, press ‘q’ to exit.

• finally, on the Internet: the right place for getting help is Usenet, like <news:comp.os.linux.setup>. Please don’t email me for help, because I’m quite overloaded.

1.3 Conventions

Throughout this work, examples will often follow the following format: <...> is a required argument, while [...] an optional one. Example:

$ tar -tf <file.tar> [> redir_file]

file.tar must be indicated, but redirection to redir_file is optional.

“RMP” means “please Read the Man Pages for further information”. I can’t stress enough how important reading the documentation is. “DOSWin” stands for “DOS/Windows”.

When the prompt of a command example is #, the command can only be performed by root.

2. For the Impatient 5

2 For the Impatient

Want to strike out? Have a look at this table:

DOS                     Linux                   Notes
------------------------------------------------------------------------------
ATTRIB (+-)attr file    chmod <mode> file       completely different
BACKUP                  tar -Mcvf device dir/   ditto
CD dirname\             cd dirname/             almost the same syntax
COPY file1 file2        cp file1 file2          ditto
DEL file                rm file                 beware - no undelete
DELTREE dirname         rm -R dirname/          ditto
DIR                     ls                      not exactly the same syntax
DIR file /S             find . -name file       completely different
EDIT file               vi file                 I think you won't like it
                        jstar file              feels like dos' edit
EDLIN file              ed file                 forget it
FORMAT                  fdformat,
                        mount, umount           quite different syntax
HELP command            man command,            same philosophy
                        info command
MD dirname              mkdir dirname/          almost the same syntax
MORE < file             less file               much better
MOVE file1 file2        mv file1 file2          ditto
NUL                     /dev/null               ditto
PRINT file              lpr file                ditto
PRN                     /dev/lp0,
                        /dev/lp1                ditto
RD dirname              rmdir dirname/          almost the same syntax
REN file1 file2         mv file1 file2          not for multiple files
RESTORE                 tar -Mxpvf device       different syntax
TYPE file               less file               much better
WIN                     startx                  poles apart!

If you need more than a table of commands, please refer to the following sections.

3 Meet bash

Good news: with Linux you type much less at the prompt, because the bash shell types for you whenever possible, and features cool line editing capabilities. To begin with, the arrow-up key recalls previous command lines; but there’s more. Pressing <TAB> completes file and directory names, so typing

$ ls /uTABloTABbTAB

is like typing

$ ls /usr/local/bin

If there were ambiguities, as typing

$ ls /uTABloTABiTAB

bash stops because it doesn’t know if you mean /usr/local/info or /usr/local/include. Supply more characters then press <TAB> again.

Other useful key presses are <ESC-BACKSPACE> that deletes a word to the left, while <ESC-D> deletes a word to the right; <ESC-F> moves the cursor one word to the right, <ESC-B> to the left; <CTRL-A> moves to the beginning of the line, <CTRL-E> to the end. The <ALT> key is equivalent to <ESC>.

Enough for now. Once you get used to these shortcuts, you’ll find the DOS prompt very annoying...

4 Files and Programs

4.1 Files: Preliminary Notions

Linux has a structure of directories and files very similar to that of DOSWin. Files have filenames that obey special rules, are stored in directories, some are executable, and among these most have command switches. Moreover, you can use wildcard characters, redirection, and piping. There are only a few minor differences:

• under DOS, file names are in the so-called 8.3 form; e.g. NOTENOUG.TXT. Under Linux we can do better. If you installed Linux using a file system like ext2 or umsdos, you can use longer filenames (up to 255 characters), and with more than one dot: for example, This_is.a.VERY_long.filename. Please note that I used both upper and lower case characters: in fact...

• upper and lower case characters in file names or commands are different. Therefore, FILENAME.tar.gz and filename.tar.gz are two different files. ls is a command, LS is a mistake;

• Windows users, beware when using long file names under Linux. If a file name contains spaces (not recommended but possible), you must enclose the file name in double quotes whenever you refer to it. For example:

     $ # the following command makes a directory called "My old files"
     $ mkdir "My old files"
     $ ls
     My old files    bin     tmp

Further, some characters shouldn’t be used: some are !*$&#.

• there are no compulsory extensions like .COM and .EXE for programs, or .BAT for batch files. Executable files are marked by an asterisk ‘*’ at the end of their name when you issue the ls -F command. For example:

     $ ls -F
     I_am_a_dir/   cindy.jpg    cjpg*   letter_to_Joe    my_1st_script*  old~

The files cjpg* and my_1st_script* are executables, that is “programs”. Under DOS, backup files end in .BAK, while under Linux they end with a tilde ‘˜ ’. Further, a file whose name starts with a dot is considered as hidden. Example: the file .I.am.a.hidden.file won’t show up after the ls command; • DOS program switches are obtained with /switch, Linux switches with -switch or –switch. Example:

dir /s becomes ls -R. Note that many DOS programs, like PKZIP or ARJ, use UNIX-style switches.

You can now jump to Section 4.4 (Translating Commands from DOS to Linux), but if I were you I’d read on.

4.2 Symbolic Links

UNIX has a type of file that doesn’t exist under DOS: the symbolic link. This can be thought of as a pointer to a file or to a directory, and can be used instead of the file or directory it points to; it’s similar to Windows shortcuts. Examples of symbolic links are /usr/X11, which points to /usr/X11R6; /dev/modem, which points

to either /dev/ttyS0 or /dev/ttyS1.

To make a symbolic link:

     $ ln -s <file_or_dir> <linkname>

Example:

     $ ln -s /usr/doc/g77/DOC g77manual.txt

Now you can refer to g77manual.txt instead of /usr/doc/g77/DOC. Links appear like this in directory listings:

     $ ls -F
     g77manual.txt@
     $ ls -l
     (several things...)           g77manual.txt -> /usr/doc/g77/DOC

4.3 Permissions and Ownership

DOS files and directories have the following attributes: A (archive), H (hidden), R (read-only), and S (system). Only H and R make sense under Linux: hidden files start with a dot, and for the R attribute, read on.

Under UNIX a file has “permissions” and an owner, who in turn belongs to a “group”. Look at this example:

     $ ls -l /bin/ls
     -rwxr-xr-x  1  root  bin  27281 Aug 15 1995 /bin/ls*

The first field contains the permissions of the file /bin/ls, which belongs to root, group bin. Leaving the remaining information aside, remember that -rwxr-xr-x means, from left to right:

- is the file type (- = ordinary file, d = directory, l = link, etc); rwx are the permissions for the file owner (read, write, execute); r-x are the permissions for the group of the file owner (read, execute); (I won’t cover the concept of group, you can survive without it as long as you’re a beginner ;-) r-x are the permissions for all other users (read, execute).

The directory /bin has permissions, too: see Section 5.2 (Directories Permissions) for further details. This is why you can’t delete the file /bin/ls unless you are root: you don’t have the permission to do so. To change a file’s permissions, the command is:

     $ chmod <whoXperm> <file>

where who is u (user, that is owner), g (group), o (other), X is either + or -, perm is r (read), w (write), or x (execute). Common examples of chmod use are the following:

     $ chmod +x file

this sets the execute permission for the file.

     $ chmod go-rw file

this removes read and write permission for everyone but the owner.

     $ chmod ugo+rwx file

this gives everyone read, write, and execute permission.

     # chmod +s file

this makes a so-called “setuid” or “suid” file—a file that everyone can execute with its owner’s privileges. Typically, you’ll come across root suid files; these are often important system files, like the X server. A shorter way to refer to permissions is with digits: rwxr-xr-x can be expressed as 755 (every letter corresponds to a bit: –- is 0, –x is 1, -w- is 2, -wx is 3...). It looks difficult, but with a bit of practice you’ll understand the concept. root, being the superuser, can change everyone’s file permissions. RMP.

4.4 Files: Translating Commands

On the left, the DOS commands; on the right, their Linux counterpart.

     ATTRIB:         chmod
     COPY:           cp
     DEL:            rm
     MOVE:           mv
     REN:            mv
     TYPE:           more, less, cat

Redirection and plumbing operators: < > >> | Wildcards: * ?

nul:  /dev/null
prn, lpt1:  /dev/lp0 or /dev/lp1; lpr

4.4.1 Examples

  DOS                                     Linux
  ---------------------------------------------------------------------
  C:\GUIDO>ATTRIB +R FILE.TXT             $ chmod 400 file.txt
  C:\GUIDO>COPY JOE.TXT JOE.DOC           $ cp joe.txt joe.doc
  C:\GUIDO>COPY *.* TOTAL                 $ cat * > total
  C:\GUIDO>COPY FRACTALS.DOC PRN          $ lpr fractals.doc
  C:\GUIDO>DEL TEMP                       $ rm temp
  C:\GUIDO>DEL *.BAK                      $ rm *~
  C:\GUIDO>MOVE PAPER.TXT TMP\            $ mv paper.txt tmp/
  C:\GUIDO>REN PAPER.TXT PAPER.ASC        $ mv paper.txt paper.asc
  C:\GUIDO>PRINT LETTER.TXT               $ lpr letter.txt
  C:\GUIDO>TYPE LETTER.TXT                $ more letter.txt
  C:\GUIDO>TYPE LETTER.TXT                $ less letter.txt
  C:\GUIDO>TYPE LETTER.TXT > NUL          $ cat letter.txt > /dev/null
          n/a                             $ more *.txt *.asc
          n/a                             $ cat section*.txt | less

Notes:

• * is smarter under Linux: * matches all files except the hidden ones; .* matches all hidden files (but also the current directory ‘.’ and parent directory ‘..’: beware!); *.* matches only those that have a ‘.’ in the middle or that end with a dot; p*r matches both ‘peter’ and ‘piper’; *c* matches both ‘picked’ and ‘peck’;

• when using more, press <SPACE> to read through the file, ‘q’ to exit. less is more intuitive and lets you use the arrow keys;

• there is no UNDELETE, so think twice before deleting anything;

• in addition to DOS’ < > >>, Linux has 2> to redirect error messages (stderr); moreover, 2>&1 redirects stderr to stdout, while 1>&2 redirects stdout to stderr;

• Linux has another wildcard: the []. Usage: [abc]* matches files starting with a, b, c; *[I-N1-3] matches files ending with I, J, K, L, M, N, 1, 2, 3;

• lpr <file> prints a file in background. To check the status of the print queue, use lpq; to remove a file from the print queue, use lprm;

• there is no DOS-like RENAME; that is, mv *.xxx *.yyy won’t work. A REN-like command is available

  on <ftp://sunsite.unc.edu/pub/Linux/utils/file>;

• use cp -i and mv -i to be warned when a file is going to be overwritten.

4.5 Running Programs: Multitasking and Sessions

To run a program, type its name as you would do under DOS. If the directory (Section 5 (Using Directories)) where the program is stored is included in the PATH (Section 8.1 (System Initialisation Files)), the program will start. Exception: unlike DOS, under Linux a program located in the current directory won’t run unless the directory is included in the PATH. Escamotage: being prog your program, type ./prog.

This is what the typical command line looks like:

  $ command [-s1 [-s2] ... [-sn]] [par1 [par2] ... [parn]] [< input] [> output]

where -s1, ..., -sn are the program switches, par1, ..., parn are the program parameters. You can issue several commands on the command line:

  $ command1 ; command2 ; ... ; commandn

That’s all about running programs, but it’s easy to go a step beyond. One of the main reasons for using Linux is that it is a multitasking os—it can run several programs (from now on, processes) at the same time. You can launch processes in background and continue working straight away. Moreover, Linux lets you have several sessions: it’s like having many computers to work on at once!

• To switch to session 1..6 on the virtual consoles, press <ALT-F1> ... <ALT-F6>

• To start a new session in the same v.c. without leaving the current one, type su - <loginname>. Example: su - root. This is useful, for instance, when you need to perform a task that only root can do.

• To end a session, type exit. If there are stopped jobs (see later), you’ll be warned.

• To launch a process in background, add an ampersand ’&’ at the end of the command line:

       $ progname [-switches] [parameters] [< input] [> output] &
       [1] 123

the shell identifies the process with a job number (e.g. [1]; see below), and with a PID (Process Identification Number; 123 in our example).

• To see how many processes there are, type ps ax. This will output a list of currently running processes.

• To kill (terminate) a process, type kill <PID>. You may need to kill a process when you don’t know how to quit it the right way.... Unless you’re root, you can’t kill other people’s processes. Sometimes, a process will only be killed by kill -SIGKILL <PID>. In addition, the shell allows you to stop or temporarily suspend a process, send a process to background, and bring a process from background to foreground. In this context, processes are called “jobs”.

• To see how many jobs there are, type jobs. Here the jobs are identified by their job number, not by their PID.

• To stop a process running in foreground, press <CTRL-C> (it won’t always work).

• To suspend a process running in foreground, press <CTRL-Z> (ditto).

• To send a suspended process into background, type bg <%job> (it becomes a job).

• To bring a job to foreground, type fg <%job>. To bring to foreground the last job sent to background, simply type fg.

• To kill a job, type kill <%job> where <job> may be 1, 2, 3,...

Using these commands you can format a disk, zip a bunch of files, compile a program, and unzip an archive all at the same time, and still have the prompt at your disposal. Try this with Windows, just to see the difference in performance (if it doesn’t crash, of course).

4.6 Running Programs on Remote Computers

To run a program on a remote machine whose name is remote.machine.edu:

     $ telnet remote.machine.edu

After logging in, start your favourite program. Needless to say, you must have a shell account on the remote machine.

If you have X11, you can even run an X application on a remote computer, displaying it on your X screen. Let remote.machine.edu be the remote X computer and let local.linux.box be your Linux machine. To run from local.linux.box an X program that resides on remote.machine.edu, do the following:

• fire up X11, start an xterm or equivalent terminal emulator, then type:

          $ xhost +remote.machine.edu
          $ telnet remote.machine.edu

• after logging in, type:

          remote:$ DISPLAY=local.linux.box:0.0
          remote:$ progname &
     (instead of DISPLAY..., you may have to write: setenv DISPLAY local.linux.box:0.0. It depends

on the remote shell.)

Et voila! Now progname will start on remote.machine.edu and will be displayed on your machine. Don’t try this over the modem though, for it’s too slow to be usable. Moreover, this is a crude and insecure method: please read the “Remote X Apps mini-HOWTO” at <http://sunsite.unc.edu/LDP/HOWTO/mini/

Remote-X-Apps>.

5 Using Directories

5.1 Directories: Preliminary Notions

We have seen the differences between files under DOSWin and Linux. As for directories, under DOSWin the root directory is \, under Linux it is /. Similarly, nested directories are separated by \ under DOSWin, by / under Linux. Example of file paths:

     DOS:    C:\PAPERS\GEOLOGY\MID_EOC.TEX
     Linux:  /home/guido/papers/geology/middle_eocene.tex

5. Using Directories 12

As usual, .. is the parent directory and . is the current directory. Remember that the system won’t let you cd, rd, or md everywhere you want. Each user has his or her stuff in a directory called ‘home’, given by the system administrator; for instance, on my PC my home dir is /home/guido.

5.2 Directories Permissions

Directories, too, have permissions. What we have seen in Section 4.3 (Permissions and Ownership) applies to directories as well (user, group, and other). For a directory, rx means you can cd to that directory, and w means that you can delete a file in the directory (according to the file’s permissions, of course), or the directory itself.

For example, to prevent other users from snooping in /home/guido/text:

$ chmod o-rwx /home/guido/text

5.3 Directories: Translating Commands

DIR:            ls, find, du
CD:             cd, pwd
MD:             mkdir
RD:             rmdir
DELTREE:        rm -rf
MOVE:           mv

5.3.1 Examples

DOS                                     Linux
---------------------------------------------------------------------
C:\GUIDO>DIR                            $ ls
C:\GUIDO>DIR FILE.TXT                   $ ls file.txt
C:\GUIDO>DIR *.H *.C                    $ ls *.h *.c
C:\GUIDO>DIR/P                          $ ls | more
C:\GUIDO>DIR/A                          $ ls -l
C:\GUIDO>DIR *.TMP /S                   $ find / -name "*.tmp"
C:\GUIDO>CD                             $ pwd
        n/a - see note                  $ cd
        ditto                           $ cd ~
        ditto                           $ cd ~/temp
C:\GUIDO>CD \OTHER                      $ cd /other
C:\GUIDO>CD ..\TEMP\TRASH               $ cd ../temp/trash
C:\GUIDO>MD NEWPROGS                    $ mkdir newprogs
C:\GUIDO>MOVE PROG ..                   $ mv prog ..
C:\GUIDO>MD \PROGS\TURBO                $ mkdir /progs/turbo
C:\GUIDO>DELTREE TEMP\TRASH             $ rm -rf temp/trash
C:\GUIDO>RD NEWPROGS                    $ rmdir newprogs
C:\GUIDO>RD \PROGS\TURBO                $ rmdir /progs/turbo

Notes:

• when using rmdir, the directory to remove must be empty. To delete a directory and all of its contents, use rm -rf (at your own risk).

• the character ‘˜ ’ is a shortcut for the name of your home directory. The commands cd or cd ˜ will take you to your home directory from wherever you are; the command cd ˜ /tmp will take you to

     /home/your_home/tmp.

• cd - “undoes” the last cd.

6 Floppies, Hard Disks, and the Like

There are two ways to manage devices under Linux: the DOS way and the UNIX way. Take your pick.

6.1 Managing Devices the DOS Way

Most Linux distributions include the Mtools suite, a set of commands that are perfectly equivalent to their DOS counterpart, but start with an ‘m’: i.e., mformat, mdir, mdel, mmd, and so on. They can even preserve long file names, but not file permissions. If you configure Mtools editing a file called /etc/mtools.conf (a sample is provided in the distribution), you can also access the DOS/Win partition, the CD–ROM, and the Zip drive. To format a fresh disk though, the mformat command won’t do. As root, you’ll have to issue this

command beforehand: fdformat /dev/fd0H1440.

You can’t access files on the floppy with a command like, say, less a:file.txt! This is the disadvantage of the DOS way of accessing disks.

6.2 Managing Devices the UNIX Way

UNIX has a different way to handle devices. There are no separate volumes like A: or C:; a disk, be it a floppy or whatever, becomes part of the local file system through an operation called “mounting”. When you’re done using the disk, before extracting it you must “unmount” it.

Physically formatting a disk is one thing, making a file system on it is another. The DOS command FORMAT A: does both things, but under Linux there are separate commands. To format a floppy, see above; to create a file system:

     # mkfs -t ext2 -c /dev/fd0H1440

You can use dos, vfat (recommended) or other formats instead of ext2. Once the disk is prepared, mount it with the command

     # mount -t ext2 /dev/fd0 /mnt

specifying the right file system if you don’t use ext2. Now you can address the files in the floppy using /mnt instead of A: or B:. Examples:

     DOS                                     Linux
     ---------------------------------------------------------------------

     C:\GUIDO>DIR A:                         $ ls /mnt
     C:\GUIDO>COPY A:*.*                     $ cp /mnt/* .
     C:\GUIDO>COPY *.ZIP A:                  $ cp *.zip /mnt
     C:\GUIDO>EDIT A:FILE.TXT                $ jstar /mnt/file.txt
     C:\GUIDO>A:                             $ cd /mnt
     A:> _/mnt/$ _

When you’ve finished, before extracting the disk you must unmount it with the command

     # umount /mnt

Obviously, you have to fdformat and mkfs only unformatted disks, not previously used ones. If you want to use the drive B:, refer to fd1H1440 and fd1 instead of fd0H1440 and fd0 in the examples above.

Needless to say, what applies to floppies also applies to other devices; for instance, you may want to mount another hard disk or a CD–ROM drive. Here’s how to mount the CD–ROM:

     # mount -t iso9660 /dev/cdrom /mnt

This was the “official” way to mount your disks, but there’s a trick in store. Since it’s a bit of a nuisance having to be root to mount a floppy or a CD–ROM, every user can be allowed to mount them this way:

• as root, do the following:

          # mkdir /mnt/floppy ; mkdir /mnt/cdrom
          # chmod 777 /mnt/floppy /mnt/cd*
          # # make sure that the CD-ROM device is right
          # chmod 666 /dev/hdb ; chmod 666 /dev/fd*

• add in /etc/fstab the following lines:

          /dev/cdrom      /mnt/cdrom  iso9660 ro,user,noauto          0       0
          /dev/fd0        /mnt/floppy vfat    user,noauto             0       0

Now, to mount a DOS floppy and a CD–ROM:

     $ mount /mnt/floppy
     $ mount /mnt/cdrom

/mnt/floppy and /mnt/cdrom can now be accessed by every user. Remember that allowing everyone to mount disks this way is a gaping security hole, if you care.

Two useful commands are df, which gives information on the mounted file systems, and du dirname which reports the disk space consumed by the directory.

6.3 Backing Up

There are several packages to help you, but the very least you can do for a multi-volume backup is (as root):

# tar -M -cvf /dev/fd0H1440 dir_to_backup/

Make sure to have a formatted floppy in the drive, and several more ready. To restore your stuff, insert the first floppy in the drive and do:

# tar -M -xpvf /dev/fd0H1440

7 What About Windows?

The “equivalent” of Windows is the graphic system X Window System. Unlike Windows or the Mac, X11 wasn’t designed for ease of use or to look good, but just to provide graphic facilities to UNIX workstations. These are the main differences:

• while Windows looks and feels the same all over the world, X11 does not: it’s much more configurable. X11’s overall look is given by a key component called “window manager”, of which you have a wide choice: fvwm, basic but nice and memory efficient, fvwm2-95, Afterstep, WindowMaker, Enlightenment, and many more. The w.m. is usually invoked from .xinitrc;

• your w.m. can be configured so as a window acts as under, er, Windows: you click on it and it comes to foreground. Another possibility is that it comes to foreground when the mouse moves over it (“focus”). Also, the placement of windows on the screen can be automatic or interactive: if a strange frame appears instead of your program, left click where you want it to appear;

• most features can be tailored editing one or more configuration files. Read the docs of your w.m.: the configuration file can be .fvwmrc, .fvwm2rc95, .steprc, etc. A sample configuration file is typically

found in /etc/X11/window-manager-name/system.window-manager-name;

• X11 applications are written using special libraries (“widget sets”); as several are available, applications look different. The most basic ones are those that use the Athena widgets (2–D look; xdvi, xman, xcalc); others use Motif (netscape), others still use Tcl/Tk, XForms, Qt, Gtk, and what have you. Nearly all of these libraries provide roughly the same look and feel as Windows;

• the feel, unfortunately, can be incoherent. For instance, if you select a line of text using the mouse and press <BACKSPACE>, you’d expect the line to disappear, right? This doesn’t work with Athena– based apps, but it does with other widget sets;

• how scrollbars and resizing work depends on the w.m. and the widget set. Tip: if you find that the scrollbars don’t behave as you would expect, try using the central button or the two buttons together to move them;

• applications don’t have an icon by default, but they can have many. Most w.m. feature a menu you recall by clicking on the desktop (“root window”); needless to say, the menu can be tailored. To change the root window appearance, use xsetroot or xloadimage; • the clipboard can only contain text, and behaves strange. Once you’ve selected text, it’s already copied

to the clipboard: move elsewhere and press the central button to paste it. There’s an application, xclipboard, that provides for multiple clipboard buffers;

• drag and drop is an option, and is only available if you use X11 applications and/or w.m. that support it.

To save memory, one should use applications that use the same libraries, but this is difficult to do in practice.

There are projects that aim at making X11 look and behave as coherently as Windows. Gnome, <http: //www.gnome.org> and KDE, <http://www.kde.org>, are awesome. Give them a try: you won’t regret your Windows desktop anymore.

8 Tailoring the System

8.1 System Initialisation Files

Two important files under DOS are AUTOEXEC.BAT and CONFIG.SYS, which are used at boot time to initialise the system, set some environment variables like PATH and FILES, and possibly launch a program or batch file. Under Linux there are lots of initialisation files, some of which you had better not tamper with until you know exactly what you are doing. I’ll tell you what the most important are, anyway:

     FILES                                   NOTES
     /etc/inittab                            don't touch for now!
     /etc/rc.d/*                             ditto

If all you need is setting the PATH and other environment variables, or you want to change the login messages or automatically launch a program after the login, have a look at the following files:

     FILES                                   NOTES
     /etc/issue                              sets pre-login message
     /etc/motd                               sets post-login message
     /etc/profile                            sets $PATH and other variables, etc.
     /etc/bashrc                             sets aliases and functions, etc.
     /home/your_home/.bashrc                 sets your aliases + functions
     /home/your_home/.bash_profile   or
     /home/your_home/.profile                sets environment + starts your progs

If the latter file exists (note that it is a hidden file), it will be read after the login, and the commands therein will be executed.

Example—look at this .bash_profile:

# I am a comment
echo Environment:
printenv | less   # equivalent of command SET under DOS

alias d='ls -l'   # easy to understand what an alias is
alias up='cd ..'
echo "I remind you that the path is "$PATH
echo "Today is `date`"  # use the output of the command 'date'
echo "Have a good day, "$LOGNAME
# The following is a "shell function"
ctgz() # List the contents of a .tar.gz archive.
{
  for file in $*
  do
    gzip -dc ${file} | tar tf -
  done
}
# end of .profile

$PATH and $LOGNAME, you guessed right, are environment variables. There are many others to play with; for instance, RMP for apps like less or bash.

Putting this line in your /etc/profile will provide the rough equivalent of PROMPT $P$G:

     export PS1="\w\\$ "

8.2 Program Initialisation Files

Under Linux, virtually everything can be tailored to your needs. Most programs have one or more initialisation files you can fiddle with, often as a .prognamerc in your home dir. The first ones you’ll want to modify are:

• .inputrc: used by bash to define key bindings;

• .xinitrc: used by startx to initialise X Window System;

• .fvwmrc: used by the window manager fvwm.

   • .joerc, .jstarrc: used by the editor joe;

• .jedrc: used by the editor jed;

• .pinerc: used by the mail reader pine;

• .Xdefault: used by many X programs.

For all of these and the others you’ll come across sooner or later, RMP. Perhaps I could interest you in the

Configuration HOWTO, <http://sunsite.unc.edu/mdw/HOWTO/Config-HOWTO.html>?

9 Networking: Concepts

Not only is “Dialup Networking” available under Linux, it’s also more stable and quicker. The name of the game is “PPP”, the protocol employed for connecting to the Internet using modems. You need a kernel that supports PPP and a tool that dials and makes the connection.

To retrieve your mail from the ISP’s server you need a tool called “email fetcher” that uses the POP protocol; when the mail is fetched it will appear as though it had been directly delivered to your Linux box. You’ll then use a MUA (Mail User Agent) like pine, mutt, elm or many others to manage it.

While under Windows the dialer is automatically invoked when you launch an Internet application, under Linux the path is the other way round: you dial first, then launch the application. A thing called diald provides the usual behaviour. Installing and configuring dialup networking used to be one of the most difficult things to do under Linux, but not anymore: please consult the Configuration HOWTO.

Finally, a word about “Network neighborhood”: you can make your Linux workstation appear as Windows NT/9x in a local network of Windows machines! The magic word is Samba: not the lively Brazilian dance, but an implementation of the SMB protocol for Linux. Go to <http://samba.anu.edu.au/samba>.

10 A Bit of Programming

10.1 Shell Scripts: .BAT Files on Steroids

If you used .BAT files to create shortcuts of long command lines (I did a lot), this goal can be attained by inserting appropriate alias lines (see example above) in profile or .bash_profile. But if your .BATs were more complicated, then you’ll love the scripting language made available by the shell: it’s as powerful as good ol’ QBasic, if not more. It has variables, structures like while, for, case, if... then... else, and lots of other features: it can be a good alternative to a “real” programming language.

To write a script—the equivalent of a .BAT file under DOS—all you have to do is write a standard ASCII file containing the instructions, save it, then make it executable with the command chmod +x <scriptfile>. To execute it, type its name.

A word of warning. The system editor is called vi, and in my experience most new users find it very difficult to use. I’m not going to explain how to use it; please consult Matt Welsh’s book or search for a tutorial on the net. Suffice it here to say that:

• to insert some text, type i then your text;

• to delete characters, type <ESC> then x;

• to quit vi whithout saving, type <ESC> then :q!

• to save and quit, type <ESC> then :wq.

A good beginner editor is joe: invoking it by typing jstar you’ll get the same key bindings as the DOSWin editor. jed in WordStar or IDE mode is even better. Please consult Section 11.4 (Where to Find Applications) to see where to get these editors.

Writing scripts under bash is such a vast subject it would require a book by itself, and I will not delve into the topic any further. I’ll just give you an example of shell script, from which you can extract some basic rules:

#!/bin/sh
# sample.sh
# I am a comment
# don't change the first line, it must be there
echo "This system is: `uname -a`" # use the output of the command
echo "My name is $0" # built-in variables
echo "You gave me the following $# parameters: "$*
echo "The first parameter is: "$1
echo -n "What's your name? " ; read your_name
echo notice the difference: "hi $your_name" # quoting with "
echo notice the difference: 'hi $your_name' # quoting with '
DIRS=0 ; FILES=0
for file in `ls .` ; do
  if [ -d ${file} ] ; then # if file is a directory
    DIRS=`expr $DIRS + 1`  # DIRS = DIRS + 1
  elif [ -f ${file} ] ; then
    FILES=`expr $FILES + 1`
  fi
  case ${file} in
    *.gif|*jpg) echo "${file}: graphic file" ;;
    *.txt|*.tex) echo "${file}: text file" ;;
    *.c|*.f|*.for) echo "${file}: source file" ;;
    *) echo "${file}: generic file" ;;
  esac
done
echo "there are ${DIRS} directories and ${FILES} files"
ls | grep "ZxY--!!!WKW"
if [ $? != 0 ] ; then # exit code of last command
  echo "ZxY--!!!WKW not found"
fi
echo "enough... type 'man bash' if you want more info."

10.2 C for Yourself

Under UNIX, the system language is C, love it or hate it. Scores of other languages (Java, FORTRAN, Pascal, Lisp, Basic, Perl, awk...) are also available.

Taken for granted that you know C, here are a couple of guidelines for those of you who have been spoilt by Turbo C++ or one of its DOS kin. Linux’s C compiler is called gcc and lacks all the bells and whistles that usually accompany its DOS counterparts: no IDE, on-line help, integrated debugger, etc. It’s just a rough command-line compiler, very powerful and efficient. To compile your standard hello.c you’ll do:

     $ gcc hello.c

which will create an executable file called a.out. To give the executable a different name, do

     $ gcc -o hola hello.c

To link a library against a program, add the switch -l<libname>. For example, to link in the math library:

     $ gcc -o mathprog mathprog.c -lm

(The -l<libname> switch forces gcc to link the library /usr/lib/lib<libname>.so; so -lm links

/usr/lib/libm.so).

So far, so good. But when your prog is made of several source files, you’ll need to use the utility make. Let’s suppose you have written an expression parser: its source file is called parser.c and #includes two header files, parser.h and xy.h. Then you want to use the routines in parser.c in a program, say, calc.c, which in turn #includes parser.h. What a mess! What do you have to do to compile calc.c?

You’ll have to write a so-called Makefile, which teaches the compiler the dependencies between sources and objects files. In our example:

# This is Makefile, used to compile calc.c
# Press the <TAB> key where indicated!
calc: calc.o parser.o
<TAB>gcc -o calc calc.o parser.o -lm
# calc depends on two object files: calc.o and parser.o
calc.o: calc.c parser.h
<TAB>gcc -c calc.c
# calc.o depends on two source files
parser.o:  parser.c parser.h xy.h
<TAB>gcc -c parser.c
# parser.o depends on three source files
# end of Makefile.

Save this file as Makefile and type make to compile your program; alternatively, save it as calc.mak and type make -f calc.mak, and of course RMP. You can invoke some help about the C functions, that are covered by man pages, section 3; for example,

     $ man 3 printf

To debug your programs, use gdb. info gdb to learn how to use it.

There are lots of libraries available; among the first you’ll want to use are ncurses (textmode effects), and svgalib (console graphics). If you feel brave enough to tackle X11 programming (it’s not that difficult), there are several libraries that make writing X11 programs a breeze. Have a look at <http://www.xnet. com/~blatura/linapp6.html>, bearing in mind that Gtk is becoming a Linux standard.

Many editors can act as an IDE; emacs and jed, for instance, also feature syntax highlighting, automatic indent, and so on. Alternatively, get the package rhide from <ftp://sunsite.unc.edu:/pub/Linux/ devel/debuggers/>. It’s a Borland IDE clone, and chances are that you’ll like it.

11 The Remaining 1%

Much more than 1%, actually...

11.1 Using tar and gzip

Under UNIX there are some widely used applications to archive and compress files. tar is used to make archives—it’s like PKZIP or Winzipbut it doesn’t compress, it only archives. To make a new archive:

     $ tar -cvf <archive_name.tar> <file> [file...]

To extract files from an archive:

     $ tar -xpvf <archive_name.tar> [file...]

To list the contents of an archive:

     $ tar -tf <archive_name.tar> | less

You can compress files using compress, which is obsolete and shouldn’t be used any more, or gzip:

     $ compress <file>
     $ gzip <file>

that creates a compressed file with extension .Z (compress) or .gz (gzip). These programs can compress only one file at a time. To decompress:

     $ compress -d <file.Z>
     $ gzip -d <file.gz>

RMP. There are also the unarj, zip and unzip (PK??ZIP compatible) utilities. Files with extension .tar.gz or .tgz (archived with tar, then compressed with gzip) are as common in the UNIX world as .ZIP files are under DOS. Here’s how to list the contents of a .tar.gz archive:

     $ tar -ztf <file.tar.gz> | less

11.2 Installing Applications

First of all: installing packages is root’s work. Most Linux applications are distributed as a .tar.gz archive, which typically will contain a directory aptly named containing files and/or subdirectories. A good rule is to install these packages from /usr/local with the command

     # tar -zxf <archive.tar.gz>

reading then the README or INSTALL file. In most cases, the application is distributed in source, which you’ll have to compile; often, typing make then make install will suffice. If the archive contains a configure script, run it first. Obviously, you’ll need the gcc or g++ compiler.

Other archives have to be unpacked from /; this is the case with Slackware’s .tgz archives. Other archives contain the files but not a subdirectory - careful not to mess things up!. Always list the contents of the archive before installing it.

Debian and Red Hat have their own archive format; respectively, .deb and .rpm. The latter is gaining wide acceptance; to install an rpm package, type

     # rpm -i package.rpm

11.3 Tips You Can’t Do Without

Backscrolling: pressing <SHIFT + PAG UP> (the grey key) allows you to backscroll a few pages, depending on how much video memory you have.

Resetting the screen: if you happen to more or cat a binary file, your screen may end up full of garbage. To fix it, blind type reset or this sequence of characters: echo CTRL-V ESC c RETURN.

Pasting text: in console, see below; in X, click and drag to select the text in an xterm window, then click the middle button (or the two buttons together if you have a two-button mouse) to paste. There is also xclipboard (alas, only for text); don’t get confused by its very slow response.

Using the mouse: if you installed gpm, a mouse driver for the console, you can click and drag to select text, then right click to paste the selected text. It works across different VCs.

Messages from the kernel: have a look at /var/adm/messages or /var/log/messages as root to see what the kernel has to tell you, including bootup messages. The command dmesg is also handy.

11.4 Where to Find Applications

If you’re wondering whether you can replace your old and trusted DOS/Win application with a Linux one, I suggest that you browse the main Linux software repositories: <ftp://sunsite.unc.edu/pub/Linux>,

<ftp://tsx-11.mit.edu/pub/linux>, and <ftp://ftp.funet.fi/pub/Linux>. Other good starting

places are the “Linux Applications and Utilities Page” <http://www.xnet.com/~blatura/linapps.shtml>, and the “official” Linux page <http://www.linux.org>.

11.5 A Few Things You Couldn’t Do

Linux can do an awful lot of things that were cumbersome, difficult or impossible do to with DOS/Windows. Here’s a short list that may whet your appetite:

• at allows you to run programs at a specified time;

• awk is a simple yet powerful language to manipulate data files (and not only). For example, being data.dat your multi field data file,

          $ awk '$2 ~ "abc" {print $1, "\t", $4}' data.dat

prints out fields 1 and 4 of every line in data.dat whose second field contains “abc”.

• cron is useful to perform tasks periodically, at specified date and time. Type man 5 crontab.

• file <filename> tells you what filename is (ASCII text, executable, archive, etc.);

• find (see also Section 5.3 (Directories: Translating Commands)) is one of the most powerful and useful commands. It’s used to find files that match several characteristics and perform actions on them. General use of find is:

     $ find <directory> <expression>

where <expression> includes search criteria and actions. Examples:

     $ find . -type l -exec ls -l {} \;

finds all the files that are symbolic links and shows what they point to.

     $ find / -name "*.old" -ok rm {} \;

finds all the files matching the pattern and deletes them, asking for your permission first.

     $ find . -perm +111

finds all the files whose permissions match 111 (executable).

     $ find . -user root

finds all the files that belong to root. Lots of possibilities here—RMP.

• grep finds text patterns in files. For example,

     $ grep -l "geology" *.tex

lists the files *.tex that contain the word “geology”. The variant zgrep works on gzipped files. RMP;

• regular expressions are a complex but darn powerful way of performing search operations on text. For example, ˆ a[ˆ a-m]X{4,}txt$ matches a line that starts with ‘a’, followed by any character except those in the interval a-m, followed by 4 or more ‘X’, and ends in ‘txt’. You use regular expressions with advanced editors, less, and many other applications. man grep for an introduction.

• script <script_file> dumps the screen contents on script_file until you issue the command exit. Useful for debugging;

• sudo allows users to perform some of root’s tasks (e.g. formatting and mounting disks; RMP);

• uname -a gives you info about your system;

• zcat and zless are useful for browsing and piping gzipped files without decompressing them. For

example:

     $ zless textfile.gz
     $ zcat textfile.gz | lpr

• The following commands often come in handy: bc, cal, chsh, cmp, cut, fmt, head, hexdump,

nl, passwd, printf, sort, split, strings, tac, tail, tee, touch, uniq, w, wall, wc,
whereis, write, xargs, znew. RMP.

11.6 Practicing UNIX under DOS/Windows

Believe it or not, there are fine tools that provide a UNIX-like environment under DOS/Windows! One is the

Djgpp suite ( <http://www.delorie.com/djgpp>) for DOS, while Cygnus ( <http://www.cygnus.com>) is

a more complex port for Win32. Both include the same GNU development tools and utilities as Linux; you won’t get the same stability and performance, though.

If you’d like to have a taste of Linux, try out Djgpp. Download and install the following files (as of this writing,

the latest version is 2.02): djdev202.zip, bnu281b.zip, bsh1147b.zip, fil316b.zip, find41b.zip,
grep22b.zip, gwk303b.zip, lss332b.zip, shl112b.zip.. Installation instructions are provided, and
you can find assistance on <news:comp.os.msdos.djgpp>.

In particular, using bash under DOSWin is a whiff of fresh air. To configure it properly, edit the supplied file BOOT.BAT to reflect your installation, then put these files in your home directory (in the Windows partition) instead of those provided:

     # this is _bashrc
     LS_OPTIONS="-F -s --color=yes"
     alias cp='cp -i'
     alias d='ls -l'
     alias l=less
     alias ls="ls $LS_OPTIONS"
     alias mv='mv -i'
     alias rm='rm -i'
     alias u='cd ..'
     # this is _bprof
     if [ -f ~/_bashrc ]; then
       . ~/_bashrc
     fi
     PS1='\w\$ '
     PS2='> '
     CDPATH="$CDPATH:~"
     # stuff for less(1)
     LESS="-M-Q"                     # long prompt, silent
     LESSEDIT="%E ?lt+%lt. %f"       # edit top line
     VISUAL="jed"                    # editor
     LESSCHARSET=latin1              # visualise accented letters
     export PS1 PS2 CDPATH LS_OPTIONS LESS LESSEDIT LESSOPEN VISUAL LESSCHARSET

11.7 Common Extensions and Related Programs

You may come across scores of file extensions. Excluding the more exotic ones (i.e. fonts, etc.), here’s a list of who’s what:

   • 1 ...  8: man pages. Read them with groff -Tascii -man <file.1>.

• arj: archive made with arj.

• dvi: output file produced by TeX (see below). xdvi to visualise it; dvips to turn it into a PostScript .ps file.

• gz: archive made with gzip.

• info: info file (sort of alternative to man pages). Get info.

• lsm: Linux Software Map file. It’s a plain ASCII file containing the description of a package.

• ps: PostScript file. To visualise or print it get gs and, optionally, ghostview or gv.

• rpm: Red Hat package. You can install it on any system using the package manager rpm.

• taz, tar.Z: archive made with tar and compressed with compress.

• tgz, tar.gz: archive made with tar and compressed with gzip.

• tex: text file to submit to TeX, a powerful typesetting system. Get the package tex, available in many distributions.

• texi: texinfo file, can produce both TeX and info files (cp. info). Get texinfo.

• xbm, xpm, xwd: graphic file.

• Z: archive made with compress.

11.8 Converting Files

If you need to exchange text files between DOS/Win and Linux, be aware of the “end of line” problem. Under DOS, each line of text ends with CR/LF (that is, ASCII 13 + ASCII 10), with LF under Linux. If you edit a DOS text file under Linux, each line will likely end with a strange–looking ‘M’ character; a Linux text file under DOS will appear as a kilometric single line with no paragraphs. There are a couple of tools, dos2unix and unix2dos, to convert the files.

If your text–only files contain accented characters, make sure they are made under Windows (with Notepad) and not under plain DOS; otherwise, all accented characters will be screwed up.

To convert Word or WordPerfect files to plain text, the matter is a bit trickier but possible. You’ll need one of the tools that can be found on the CTAN sites; one is <ftp://ftp.tex.ac.uk>. Get the program word2x from the directory /pub/tex/tools/, or try one the programs available in directory /pub/tex/support/. In particular, word2x converts Word 6 files, while for Word 97 files you’ll need mswordview ( <http://www.

csn.ul.ie/~caolan/docs/MSWordView.html>) that turns them to HTML.

11.9 Free Office Suites

If converting files is not enough, you have the choice of sinking your teeth into a (free!) Microsoft Office–like package.

The StarOffice suite is free for private use. It’s big, somewhat slow, but very good anyway: it offers a lot of functionality not found in Microsoft Office. It can also read and write Word and Excel files, although the conversion isn’t always perfect. Home page: <http://www.stardivision.com>.

Another good package is Corel WordPerfect, a free edition of which is available for download. Should I say

more? Go fetch it: <http://www.corel.com>.

12 The End, for Now

Congratulations! You have now grasped a little bit of UNIX and are ready to start working. Remember that your knowledge of the system is still limited, and that you are expected to do more practice with Linux to use it comfortably. But if all you had to do was get a bunch of applications and start working with them, what I included here is enough.

I’m sure you’ll enjoy using Linux and will keep learning more about it—everybody does. I bet, too, that you’ll never go back to DOSWin! I hope I made myself understood and did a good service to my 3 or 4 readers.

12.1 Copyright

Unless otherwise stated, Linux HOWTO documents are copyrighted by their respective authors. Linux HOWTO documents may be reproduced and distributed in whole or in part, in any medium physical or electronic, as long as this copyright notice is retained on all copies. Commercial redistribution is allowed and encouraged; however, the author would like to be notified of any such distributions.

All translations, derivative works, or aggregate works incorporating any Linux HOWTO documents must be covered under this copyright notice. That is, you may not produce a derivative work from a HOWTO and impose additional restrictions on its distribution. Exceptions to these rules may be granted under certain conditions; please contact the Linux HOWTO coordinator at the address given below.

In short, we wish to promote dissemination of this information through as many channels as possible. However, we do wish to retain copyright on the HOWTO documents, and would like to be notified of any plans to redistribute the HOWTOs.

If you have questions, please contact Tim Bynum, the Linux HOWTO coordinator, at <linux-howto@

sunsite.unc.edu> via email.

12.2 Disclaimer

“From DOS to Linux HOWTO” was written by Guido Gonzato, <REMOVE_MEguido@ibogeo.df.unibo.it>. (Remove “REMOVE_ME”.) Many thanks to Matt Welsh, the author of “Linux Installation and Getting Started”, to Ian Jackson, the author of “Linux frequently asked questions with answers”, to Giuseppe Zanetti, the author of “Linux”, to all the folks who emailed me suggestions, and especially to Linus Torvalds and GNU who gave us Linux.

This document is provided “as is”. I put great effort into writing it as accurately as I could, but you use the information contained in it at your own risk. In no event shall I be liable for any damages resulting from the use of this work.

Feedback is welcome. For any requests, suggestions, flames, etc., feel free to contact me.

Enjoy Linux and life,

Guido =8-)